1.Find fxy(x,y) if f(x,y)=(x^5+y^4)^6.
2. Find Cxy(x,y) if C(x,y)=6x^2-3xy-7y^2+2x-4y-3


Find (,,(Xy) if f(x,y)= (x + y) fxy(x,y) =
Find Cxy(x,y) if C(x,y) = 6x² + 3xy – 7y2 + 2x - 4y - 3. Cxy(x,y)=0
Answer #1

(1)

Gine, f (a,y)= 5 af = ay f (x,y) = (x + ayu) 6 fy (2,4)= 6 ( 2 + yh). (443). f (x,y) = 2453 ( x +ya) 3 (f: (1.1)) = 24 go {36

(2)

Given, - 3, Әe ( х x is treated ССма) 6х++ 38 -13° + 9+ - - 3 с ( 3) = Эх - 148 - 1 2 ( ча) - 3 as constant). ә4 -с 3, Coxy (